Moody Corporation uses a job-order costing system with a plantwide predetermined overhead rate based on machine-hours. At the beginning of the year, the company made the following estimates:

Machine-hours required to support estimated production 151,000
Fixed manufacturing overhead cost $ 657,000
Variable manufacturing overhead cost per machine-hour $ 4.30

Required:

1. Compute the plantwide predetermined overhead rate.

Complete this question by entering your answers in the tabs below.

  • Required 1
  • Required 2
  • Required 3
  • Required 4

Compute the plantwide predetermined overhead rate. (Round your answer to 2 decimal places.)

Predetermined overhead rate per MH   


2. During the year, Job 400 was started and completed. The following information was available with respect to this job:

Direct materials $ 350
Direct labor cost $ 290
Machine-hours used 38

Compute the total manufacturing cost assigned to Job 400. (Round your intermediate calculations to 2 decimal places and your final answer to the nearest dollar amount.)

Total manufacturing cost
  


Compute the total manufacturing cost assigned to Job 400.

3. If Job 400 includes 50 units, what is the unit product cost for this job?

If Job 400 includes 50 units, what is the unit product cost for this job? (Round your intermediate calculations and final answer to the nearest whole dollar.)

Unit product cost

4. If Moody uses a markup percentage of 120% of its total manufacturing cost, then what selling price per unit would it have established for Job 400?

If Moody uses a markup percentage of 120% of its total manufacturing cost, then what selling price per unit would it have established for Job 400? (Round your intermediate calculations and final answer to the nearest whole dollar.)

Selling price per unit
Answer #1

1.

The plantwide predetermined overhead rate :
The plantwide predetermined overhead rate = Variable Overhead Rate + Fixed Overhead Rat
Fixed Overhead Rate = Fixed Manufacturing Cost / Estimated Machine Hours
Fixed Overhead Rate =$657,000 / 151,000 = $4.35
Predetermined overhead rate ($4.30 + $4.35) $ 8.65 per MH   

2.

Job 400
Total Manufacturing Cost
Direct Material $      350.00 350
Direct Labor $      290.00 290
Manufacturing Overhead
Variable: $      163.40 =4.30*38
Fixed $      165.30 =4.35*38
Total Manufacturing Cost $      968.70

3.

Unit product cost = Total Manufacturing Cost / Total Units
Unit product cost = $968.70 / 50
Unit Product Cost = 19.37 per unit  

4.

Selling price per unit
Total Manufacturing Cost per unit $        19.37
Add: Markup Percentage (120%) $        23.24
Selling price per unit $        42.61